# WaterVG: Waterway Visual Grounding based on Text-Guided Vision and mmWave Radar

### v1

@@ Abstract.
The perception of waterways based on human intent holds significant importance for autonomous navigation and operations of Unmanned Surface Vehicles (USVs) in water environments. Inspired by visual grounding, in this paper, we introduce WaterVG, the first visual grounding dataset designed for USV-based waterway perception based on human intention prompts. WaterVG encompasses prompts describing multiple targets, with annotations at the instance level including bounding boxes and masks. Notably, WaterVG includes 11,568 samples with 34,950 referred targets, which integrates both visual and radar characteristics captured by monocular camera and millimeter-wave (mmWave) radar, enabling a finer granularity of text prompts. Furthermore, we propose a novel multi-modal visual grounding model, Potamoi, which is a multi-modal and multi-task model based on the one-stage paradigm with a designed Phased Heterogeneous Modality Fusion (PHMF) structure, including Adaptive Radar Weighting (ARW) and Multi-Head Slim Cross Attention (MHSCA). In specific, MHSCA is a low-cost and efficient fusion module with a remarkably small parameter count and FLOPs, elegantly aligning and fusing scenario context information captured by two sensors with linguistic features, which can effectively address tasks of referring expression comprehension and segmentation based on fine-grained prompts. Comprehensive experiments and evaluations have been conducted on WaterVG, where our Potamoi archives state-of-the-art performances compared with counterparts.
Visual grounding, Multi-modal fusion, USV-based water-surface perception, mmWave radar
This work is partially supported by the XJTLU AI University Research Centre and Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU. Also, it is partially funded by the Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004) as well as funding: XJTLU-REF-21-01-002, XJTLU- RDF-22-01-062, and XJTLU Key Program Special Fund (KSF-A-17). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).Runwei Guan, Liye Jia and Fengyufan Yang contribute equally to this work.*Corresponding email: yutaoyue@hkust-gz.edu.cn.
@@ 1. Introduction
With the rapid advancement of deep learning technologies, intelligent perception in water-surface environments has witnessed substantial progress in various domains (Yao et al., 2023a, b), including autonomous driving of USVs (Yao et al., 2023c), maritime intelligent transportation (Guo et al., 2023), etc. Specifically, current advancements in water-surface intelligent perception primarily encompass instance-level detection (Cheng et al., 2021c), segmentation(Guan et al., 2023c, b), and tracking of waterborne targets (Zhang et al., 2021), detection of drivable areas (Guan et al., 2024), waterline detection (Guan et al., 2023d), Simultaneous Localization And Mapping (SLAM) of waterways (Cheng et al., 2021a), and related facets. Based on the aforementioned investigation, our findings reveal that current research predominantly focuses on sensor-based automatic perception of water surface environments.
However, there is a paucity of research on human-guided specific target perception through natural language, which belongs to the domain of visual grounding. In the context of visual grounding in waterways, humans can convey target appearance, position, distance and related features through descriptions, facilitating more intuitive and efficient guidance for USVs. This leads to enhanced precision in navigation and localization, thereby holding significant implications for waterway intelligent traffic systems, search and rescue operations, and aquatic environment conservation. Consequently, our focus is centered on the domain of visual grounding, encompassing two primary tasks: Referring Expression Comprehension (REC) (Qiao et al., 2020) and Referring Expression Segmentation (RES) (Chrupała, 2022). These tasks involve presenting a prompt that describes one or more targets, and the visual grounding model is tasked with predicting bounding boxes and masks of the objects in the image that most closely match the description provided in the prompt.
Building upon the aforementioned considerations, we contemplate the establishment of an optimal visual grounding dataset tailored for robust perception and monitoring in waterway environments. Firstly, pure visual sensors are susceptible to interference in adverse weather and lighting (Cheng et al., 2021c; Yao et al., 2023a, c). Secondly, visual sensors can only capture qualitative features of targets in terms of appearance or position. but traffic supervisors often need quantitative distance and motion information about waterway targets. Therefore, when one desires to ascertain whether there is a vessel approaching behind the fog or whether obstacles exist within a specific distance, visual sensors are unable to precisely locate targets based on human intent. To enable target referring flexibly, we introduce 4D mmWave radar to complement the visual sensor, which as an all-weather perception sensor, can detect targets in the form of point clouds, offering higher resolution compared to conventional mmWave radars (Guan et al., 2023b, d; Cheng et al., 2021b). Concurrently, while visual inputs provide appearance features, including color, size, texture and shape, radar can furnish target reflection characteristics (Yao et al., 2023b), including distance, velocity, reflected power, etc.
@@
Figure 2. Samples in WaterVG, containing referred targets with prompts under various situations. The annotations include bounding boxes and masks, supplemented by radar point clouds (red).
@@
Upon these insights, we develop the first visual grounding dataset for water-surface perception, WaterVG, based on the WaterScenes dataset (Yao et al., 2023c), which is a large-scale USV-based waterway perception dataset integrating the monocular camera and 4D mmWave radar. Detailedly, diverging from other vision-based visual grounding datasets, our WaterVG dataset has the following distinguishing features: i) Manual fine-grained prompts with a maximum of 40 tokens to describe targets not only with visual attributes like color, size, shape and texture but also incorporate features imperceptible to vision yet detectable by radar, such as numeric distance and motion characteristics. ii) Flexible referred targets with 1 to 15 per prompt. iii) Diverse prompts, containing specific descriptions, coarse and fine queries about one or more of the target’s characteristics.
Correspondingly, we propose a novel baseline model, Potamoi, adept at seamlessly integrating these three distinctly heterogeneous modalities. The crux of fusing heterogeneous modalities lies in how to align their diverse features within a latent space. Building on this, we propose an effective fusion structure called Phased Heterogeneous Modality Fusion (PHMF). PHMF performs a phased fusion of visual, radar, and linguistic features. This approach not only effectively alleviates the complexity increase and speed reduction caused by the superposition of multiple modalities with multiple branching channels but also deepens the neural network, enhancing the model’s expressive capacity. PHMF’s core lies in our proposed multi-modal fusion method called Multi-Head Slim Cross Attention (MHSCA). This module within Potamoi efficiently fuses abundant contextual information captured by two sensors with fine-grained linguistic features, achieving exemplary performance on both REC and RES tasks with lower parameter numbers and FLOPs than vanilla Multi-Head Cross Attention (MHCA) (Wu et al., 2023) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
@@
    We establish the first visual grounding dataset for USV-based waterway perception, WaterVG, which comprises 11,568 triplet input data, including images, radar point clouds, and prompts. Moreover, the ground truth encompasses two types of instance-level annotations: bounding boxes and masks. In particular, our textual prompts contain extra corresponding descriptions of target distance and motion information.
@@
    We propose a novel one-stage multi-task baseline model, Potamoi, where an effective fusion network structure called Phased Heterogeneous Modality Fusion (PHMF) aligns and fuses three modalities. PHMF includes Adaptive Radar Weighting (ARW) modules and a low-cost fusion method called Multi-Head Slim Cross Attention (MHSCA). MHSCA can fuse heterogeneous modalities efficiently for visual grounding with few parameters, low FLOPs and power than vanilla Multi-Head Cross Attention (MHCA) (Dao et al., 2022) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
@@
    Comprehensive experiments based on the WaterVG dataset encompasses various aspects such as the performance of visual grounding models under fusion and vision patterns, and multi-modal fusion methods. These experiments provide further insights and deepen our understanding of the field.
@@ 2. Related Works
2.1. Datasets of Visual Grounding
@@
As a challenging branch within multi-modal learning, visual grounding aims to locate, based on natural language query, the targets in images that best match the descriptions provided by the natural language (Zhou et al., 2023). This is further divided into REC and RES tasks. The REC task focuses on localizing targets in the form of bounding boxes, whereas RES aims to delineate target masks. The majority of current 2D visual grounding datasets (Kazemzadeh et al., 2014; Plummer et al., 2015; Yu et al., 2016; Krishna et al., 2017; Deruyttere et al., 2019; Wang et al., 2020; Chen et al., 2020; Zhan et al., 2023; Liu et al., 2023; Wu et al., 2023; He et al., 2023) primarily cater to REC tasks, relying solely on pure visual data, lacking other modalities queried by prompts. Moreover, there is a current trend towards a prompt being capable of querying multiple targets, deviating from the single-target prompt approach as seen in datasets like RefCOCOs (Yu et al., 2016), thereby aligning more flexibly with real-world scenarios. Additionally, multi-task learning stands as a significant research focus within visual grounding. In contrast to other datasets, our WaterVG supports both REC and RES tasks, accommodating natural language reference to multiple targets and facilitating multi-task learning. Notably, we provide image and radar point cloud data concurrently, enabling more precise queries based on natural language prompts.
2.2. Visual Grounding Models
As visual grounding advances, its models are primarily categorized into three major classes: two-stage (Liu et al., 2019; Yang et al., 2019b; Yu et al., 2018; Wang et al., 2019; Hong et al., 2019), one-stage (Luo et al., 2020; Chen et al., 2018; Yang et al., 2019a), and transformer-based models (Zhu et al., 2022; Du et al., 2022; Deng et al., 2021). Among them, firstly, two-stage models split the generation of Regions Of Interest (ROI) and target classification into two parts, where the first stage is responsible for matching the regions of interest with text features while the second stage is to classify the foreground and background categories of ROI based on the fused features of text and image. Therefore, two-stage models usually adopt Faster R-CNN (Girshick, 2015) as the skeleton. Secondly, one-stage models usually combine bounding box regression and classification into one step. Besides, linguistic features usually fuse with image features at the end of the image backbone or within Feature Pyramid Networks (FPN). Thirdly, transformer-based models are usually in the format of an image backbone and a text encoder followed by a transformer-based encoder and decoder, where the fusion operation happens before the transformer-based encoder. There are many different text-visual fusion methods among the aforementioned models, including cross-modal spatial attention (Zhu et al., 2022), multi-head cross attention (Wu et al., 2023), channel concatenation (Deng et al., 2021), and so on.
@@ 3. The WaterVG Dataset
Figure 3. Five-step prompt annotations based on the samples of image and projected radar point clouds.
@@ 3.1. Dataset Construction
Our WaterVG dataset is built upon the WaterScenes dataset (Yao et al., 2023c), a high-quality, multi-sensor dataset with vision and radar, specifically designed for various water-surface environments. It features high-resolution images from a monocular camera with 1920 × 1080 pixels, complemented by 4D radar point clouds, which are calibrated and time-synchronized with the monocular camera. Additionally, WaterScenes (Yao et al., 2023c) provides high-quality annotations for object detection in the form of bounding boxes, as well as instance segmentation masks. Building upon these resources, we engage two experts with extensive experience in intelligent aquatic transportation to conduct prompt annotation of our WaterVG dataset, which selects 11,568 representative samples from WaterScenes under different scenarios and lighting conditions. As illustrated in Fig. 3, the annotation process is divided into five distinct stages.
@@
(1) Select candidate targets. To facilitate efficient annotation by experts, we develop an annotation GUI client specifically for the WaterVG dataset. In this system, the radar point clouds are projected onto the camera plane using the projection matrix based on the exterior parameters of the radar and camera, as well as the intrinsic parameters of the camera, with the point clouds displayed in red. Each point cloud is accompanied by information displaying the distance of the target from the USV and its radial velocity. Correspondingly, both bounding box and mask annotations are mapped onto the image. Annotators simply need to click on the location of the target to display these two types of annotations.
@@
(2) Check the annotations. Once the bounding box and mask of a clicked object are displayed on the image, two annotators proceed to check the corresponding ground truth for accuracy. If the ground truth is verified as correct, they continue the next step. If not, the target is re-annotated to ensure the precision of the dataset.
@@
(3) Write referring prompts. Based on the images and radar point clouds, the two annotators include as many physical characteristics of the targets as possible in the annotation prompt. These encompass the target’s category, size, color, and shape features perceptible by visual means. Furthermore, if the target is detected by the radar precisely accompanied by dense projected point clouds, we use characteristics of point clouds to further describe the target’s distance and motion trends from our USV. Additionally, if multiple targets are involved, annotators will also work on expressing their spatial location relationships. Based on these, the annotators aim to diversify the prompts by incorporating various sentence structures and methods of referencing the target.
@@
(4) Enrich prompts. When facing identical scenes or targets in different samples, the limited vocabulary and subjective tone of the annotators are unavoidable. To enhance the diversity of prompt content, we utilized the existing annotated prompts and fed them into the ChatGPT-3.5-Turbo API, aiming to enrich and diversify the prompts by generating semantically equivalent yet more elaborate and varied expressions.
@@
(5) Check each other’s prompt annotations. To avoid ambiguity in text prompts, as a final step, we exchange prompts written by two experts. Each expert reviews and examines the text prompt written by the other, respecting their respective writing styles. In cases where there is ambiguity or misidentification in target references, the two experts engage in discussions and reach a consensus for necessary modifications.
Figure 4. The statistics of WaterVG. (a) waterway proportion, (b) sentence pattern proportion, (c) query type proportion, (d) prompt length per sample, (e) word cloud, (f) co-occurrence of categories, (g) referred target number per sample, (h) category distribution of referred targets.

### v2

@@ Abstract.
The perception of waterways based on human intent is significant for autonomous navigation and operations of Unmanned Surface Vehicles (USVs) in water environments. Inspired by visual grounding, we introduce WaterVG, the first visual grounding dataset designed for USV-based waterway perception based on human prompts. WaterVG encompasses prompts describing multiple targets, with annotations at the instance level including bounding boxes and masks. Notably, WaterVG includes 11,568 samples with 34,987 referred targets, whose prompts integrates both visual and radar characteristics. The pattern of text-guided two sensors equips a finer granularity of text prompts with visual and radar features of referred targets. Moreover, we propose a low-power visual grounding model, Potamoi, which is a multi-task model with a well-designed Phased Heterogeneous Modality Fusion (PHMF) mode, including Adaptive Radar Weighting (ARW) and Multi-Head Slim Cross Attention (MHSCA). Exactly, ARW extracts required radar features to fuse with vision for prompt alignment. MHSCA is an efficient fusion module with a remarkably small parameter count and FLOPs, elegantly fusing scenario context captured by two sensors with linguistic features, which performs expressively on visual grounding tasks. Comprehensive experiments and evaluations have been conducted on WaterVG, where our Potamoi archives state-of-the-art performances compared with counterparts.
Waterway visual grounding, Multi-modal fusion, Low-power model, USV-based perception, mmWave radar
submissionid: 687
@@ 1. Introduction
With the advancement of deep learning, water-surface intelligent perception has witnessed substantial progress in various domains (Yao et al., 2023a, b, c; Guo et al., 2023). Specifically, current advancements primarily encompass object detection (Cheng et al., 2021c; Guan et al., 2023d), segmentation(Guan et al., 2023c, b, 2024), tracking (Zhang et al., 2021) and Simultaneous Localization And Mapping (SLAM) (Cheng et al., 2021a). Based on the aforementioned, waterway perception faces the following challenges, firstly, waterways are often characterized by complexity, irregularity, and variability in obstacles and lighting conditions. Secondly, due to fluctuations in water levels and changes in flow velocity, many obstacles are irregular and shifting. Thirdly, vessel navigation in waterways usually does not follow prescribed routes, leading to a higher frequency of unforeseen events and increased challenges in waterway monitoring. Therefore, it is challenging for regulators or specialized USVs to locate desired waterway targets.
However, current sensor-based full-target perception may not adequately address this issue. Hence, we pioneeringly introduce visual grounding to waterway perception, utilizing natural language as prompt to locate specific targets. Here, humans can convey target exact features through descriptions, facilitating more intuitive and efficient guidance for USVs. This leads to enhanced precision in navigation and localization, thereby holding significant implications for transportation, rescue operations, and environment conservation.
@@
Figure 2. Samples in WaterVG. The annotations include bounding boxes and masks, supplemented by radar point clouds (red).
Building upon the aforementioned considerations, we contemplate the establishment of an optimal visual grounding dataset tailored for robust perception and monitoring in waterway environments. Here are our considerations, firstly, pure visual sensors are easily disturbed by variable visibility and lighting conditions (Cheng et al., 2021c; Yao et al., 2023a, c). Secondly, visual sensors can only capture qualitative features of targets in terms of appearance or position, but traffic supervisors often need quantitative distance and motion information about waterway targets. Therefore, when one desires to ascertain whether there is a vessel approaching behind the fog or whether obstacles exist within a specific distance, visual sensors are unable to precisely locate targets based on human intent. To enable target referring flexibly, we complement the visual sensor with 4D mmWave radar, as it is an all-weather sensor that can detect targets in the form of point clouds and offer higher resolution than conventional mmWave radar (Guan et al., 2023b, d; Cheng et al., 2021b). Concurrently, while visual sensors perceive appearance features, including color, size, texture and shape, radar can furnish target reflection characteristics (Yao et al., 2023b), including distance, velocity, reflected power, etc.
@@
Upon these insights, we build the first visual grounding dataset called WaterVG for waterway perception, based on the WaterScenes dataset (Yao et al., 2023c), which is a USV-based waterway perception dataset integrating the monocular camera and 4D radar. Detailedly, diverging from other vision-based visual grounding datasets, our WaterVG dataset has the following distinguishing features: i) Manual fine-grained prompts to describe targets through both visual attributes like color, size, shape, texture and features detectable by radar, such as numeric distance and motion characteristics. ii) Flexible referred targets with 1 to 15 per prompt. iii) Diverse prompts, containing specific descriptions, coarse and fine queries on single or multiple characteristics. iv) Multi-task annotations of Referring Expression Comprehension (Qiao et al., 2020) and Segmentation (Chrupała, 2022) (REC and RES).
Correspondingly, three heterogeneous modalities, RGB image, 3D point clouds and natural language will face several challenges regarding fusion for USV-based visual grounding. Firstly, radar may be interfered with clutter, causing some non-target points (ghost echoes). Hence, it is crucial to adaptively extract radar features mentioned in prompts while ignoring clutter. Secondly, referred targets vary significantly in visual and point cloud scale, so the dynamic fusion of sparse sensor features with corresponding dense language features is vital. Thirdly, fragmented feature extraction and fusion branches of multi-modal networks tend to cause increased power consumption and reduction in USV endurance time.
Therefore, we propose a low-power multi-task model called Potamoi, achieving a trade-off between accuracy and power consumption. Within this, we design an effective fusion mode called Phased Heterogeneous Modality Fusion (PHMF). PHMF performs a phased alignment and fusion of visual, radar, and linguistic features. In PHMF, Adaptive Radar Weighting (ARW) is proposed for high-quality radar feature extraction, effectively suppressing interference from clutter in space and adaptively extracting channel features of target radar characteristics contained within text prompts to the maximum extent. More importantly, PHMF’s core lies in our proposed Multi-Head Slim Cross Attention (MHSCA). MHSCA efficiently reduces redundancy in query context and fuses contextual features captured by two sensors with linguistic features, achieving exemplary performance on visual grounding with lower parameters and FLOPs than vanilla Multi-Head Cross Attention (MHCA) (Wu et al., 2023) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
@@
    The first visual grounding dataset for USV-based waterway perception, WaterVG, comprises 11,568 triplet input data, including images, radar point clouds, and prompts. Moreover, the ground truth encompasses two types of instance-level annotations: bounding boxes and masks. In particular, our textual prompts contain extra corresponding descriptions of target distance and motion features.
@@
    A low-power multi-task visual grounding model, Potamoi, where an effective fusion mode called PHMF, aligns and fuses three modalities. PHMF includes ARW for radar feature adaptive extraction and a low-cost fusion method called MHSCA, which can fuse heterogeneous modalities efficiently for visual grounding with few parameters, low FLOPs and power than vanilla MHCA and MHLCA.
@@
    Comprehensive experiments based on the WaterVG dataset encompass various aspects such as the performance of visual grounding models under fusion and vision patterns, and multi-modal fusion methods. These experiments deepen insights and understanding of the field.
Figure 3. The statistics of WaterVG, including proportion on the proportion of waterways (a), sentence patterns (b) and query types (c), and the statistics of WaterVG on referred target number (d), category distribution (e), prompt length (f) and word cloud (g).
@@ 2. Related Works
2.1. Waterway Perception Datasets
Recently, waterway perception datasets have obtained significant attention in the field of waterway robotics and autonomous navigation. Most waterway perception datasets cover instance-level water-surface obstacle detection (Kristan et al., 2015; Bovcon et al., 2018; Liu et al., 2021; Cheng et al., 2021c), segmentation (Bovcon et al., 2019; Taipalmaa et al., 2019; Zhou et al., 2021; Žust et al., 2023) and tracking (Moosbauer et al., 2019; Varga et al., 2022) for traffic monitoring, autonomous driving and water environmental protection. Furthermore, datasets for the SLAM (Miller et al., 2018; Cheng et al., 2021a) and segmentation (Yao et al., 2023c; Cheng et al., 2021a) of drivable-area and shorelines are also proposed for the planning-oriented perception of USVs. Based on the above observations, it is evident that current waterway perception datasets primarily focus on sensor-based automatic perception, lacking flexibility in human-machine interaction. Text prompts can be regarded as a way to simulate human instructions. In many scenarios, leveraging text prompts, USVs (waterborne robots) can more flexibly locate and identify specific targets, thereby achieving more efficient execution of waterway tasks. Therefore, we design and construct the first waterway visual grounding dataset, WaterVG, to promote human-guided waterway monitoring and USV-based embodied AI in waterway perception.
2.2. Visual Grounding Datasets and Models
@@
As a challenging branch within multi-modal learning, visual grounding aims to locate, based on text query, the targets in images that best match the descriptions provided by the natural language (Zhou et al., 2023). This is further divided into REC and RES tasks. The REC task focuses on localizing targets with bounding boxes, whereas RES aims to delineate target masks. The majority of current 2D visual grounding datasets (Kazemzadeh et al., 2014; Plummer et al., 2015; Yu et al., 2016; Krishna et al., 2017; Deruyttere et al., 2019; Wang et al., 2020; Chen et al., 2020; Zhan et al., 2023; Liu et al., 2023; Wu et al., 2023; He et al., 2023) primarily cater to REC tasks, relying solely on pure visual data, lacking features from other synchronous modalities queried by prompts. Moreover, there is a current trend towards a prompt capable of querying multiple targets, which aligns more flexibly with real-world scenarios, deviating from the single-target reference. Moreover, visual grounding models are primarily categorized into three classes: two-stage (Liu et al., 2019; Yang et al., 2019b; Yu et al., 2018; Wang et al., 2019; Hong et al., 2019), one-stage (Luo et al., 2020; Chen et al., 2018; Yang et al., 2019a), and transformer-based models (Zhu et al., 2022; Du et al., 2022; Deng et al., 2021). Among them, firstly, two-stage models split the generation of Regions Of Interest (ROI) and target classification into two parts, where the first stage is responsible for matching the regions of interest with text features while the second stage is to filter and refine ROI based on the fused features of text and image. Secondly, one-stage models usually combine bounding box regression and confidence prediction into one step. Besides, linguistic features usually fuse with image features at the end of the image backbone or within multi-scale fusion networks. Thirdly, transformer-based models are usually in the format of an image backbone and a text encoder followed by a transformer-based encoder and decoder, where the fusion operation usually happens before the transformer-based encoder.
@@ 3. The WaterVG Dataset
Figure 4. Five-step prompt annotations of WaterVG.
@@ 3.1. Dataset Construction
Our WaterVG dataset is built upon the WaterScenes dataset (Yao et al., 2023c), a high-quality, multi-sensor dataset with vision and radar, specifically designed for various water-surface environments. It features high-resolution images from a monocular camera with 1920 × 1080 pixels, complemented by 4D radar point clouds, which are calibrated and time-synchronized with the monocular camera. Additionally, WaterScenes (Yao et al., 2023c) provides high-quality annotations for object detection in the form of bounding boxes, as well as segmentation masks. Building upon these resources, we engage two experts with extensive experience in intelligent waterway transportation to conduct prompt annotation of our WaterVG dataset, selecting 11,568 representative samples from WaterScenes under different scenarios and lighting conditions. As illustrated in Fig. 4, the annotation process is divided into five distinct stages.
@@
(1) Select candidate targets. To facilitate efficient annotation by experts, we develop an annotation GUI client specifically for the WaterVG dataset. In this system, the radar point clouds are projected onto the camera plane using the projection matrix based on the extrinsic parameters of the radar and camera, as well as the intrinsic parameters of the camera, with the point clouds displayed in red. Each point cloud is accompanied by information displaying the distance of the target from the USV and its radial velocity. Correspondingly, both bounding box and mask annotations are mapped onto the image. Annotators simply need to click on the location of the target to display these two types of annotations.
@@
(2) Check the annotations. Once the bounding box and mask of a clicked object are displayed on the image, two annotators proceed to check the accuracy of corresponding ground truth. If the ground truth is verified as correct, they continue the next step. If not, the target is re-annotated to ensure the precision of the dataset.
@@
(3) Write referring prompts. Based on the images and radar point clouds, two annotators include as many physical characteristics of the targets as possible in the annotation prompt. These encompass the target’s category, size, color, and shape features perceptible by visual means. Furthermore, if the target is detected by the radar precisely accompanied by dense projected point clouds, we use characteristics of point clouds to further describe the target’s distance and motion trends from our USV. Additionally, if multiple targets are involved, annotators will also work on expressing their spatial location relationships. Based on these, the annotators aim to diversify the prompts by incorporating various sentence structures and methods of referencing the target.
@@
(4) Enrich prompts. When facing identical scenes or targets in different samples, the limited vocabulary and subjective tone of annotators are unavoidable. To enhance the diversity of prompt content, we utilize the existing annotated prompts and feed them into ChatGPT-3.5-Turbo API (Ouyang et al., 2022), aiming to enrich the prompts by generating semantically equivalent yet more varied expressions.
@@
(5) Check each other’s prompt annotations. To avoid ambiguity in prompts, we exchange prompts written by two experts. Each expert reviews and examines prompts written by the other, respecting respective writing styles. In cases where there is ambiguity or misidentification in target references, the two experts engage in discussions and reach a consensus for necessary modifications.